How much agarose would you need to weigh out in order to make 65ml of a 0.7% gel? Show your calculation
Answer #1

Answer :-

0.7 % means 700 mg in 1000 ml of water.

1000 ml of water = 700 mg

1 ml of water = 700/1000 = 0.7 mg of agar

So 65 ml of water will have 0.7×65 = 45.5 mg of agar.

So you need to weigh 45.5 mg of agar.